Getting a defective item when two items are randomly selected from a box of two defective and three non-defective items       


1
Expert's answer
2022-03-07T17:15:04-0500

Let A - ''getting a defective item'', then "P(A)=1-P(\neg A)"

"P(\neg A)={\frac m n}" , where n - number of options of taking two items out of five, m - number of options of taking none defective item

"P(A)=1-{\frac {{ 3\choose 2}} {{5 \choose 2}}}=1-{\frac 3 {10}}={\frac 7 {10}}"